# In-Context Learning as Nonparametric Conditional Probability Estimation: Risk Bounds and Optimality

### v1

@@ Appendix A Proof of Theorem 7
  • The expected KL risk defined as:
@@
  • The empirical risk gap is defined as:
@@ A.1 Supporting Lemmas
$\bigg{|}\mathbb{E}_{\mathcal{D}_{T}}\left[y_{k}^{(t)}\left(B\wedge\log\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{\hat{p}_{k}(\mathbf{Z}^{(t)})}\right)\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})<\frac{C_{T}}{T}\right)\right]\bigg{|}\leq\frac{C_{T}\{\log(T/C_{T})+B\}}{T},$
@@
$\bigg{|}\mathbb{E}_{\mathcal{D}_{T}}\left[y_{k}^{(t)}\left(B\wedge\log\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{\hat{p}_{k}(\mathbf{Z}^{(t)})}\right)\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})<\frac{C_{T}}{T}\right)\right]\bigg{|}$
@@
This complete the proof of this lemma.
@@
Recall that $\mathbf{p}(\mathbf{Z})=(p_{1}(\mathbf{Z}),\dots,p_{K}(\mathbf{Z}))$ and $g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})=\mathbf{y}^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\mathbf{p}(\mathbf{Z})}\right)_{\geq C_{T}/T}$, where $B\geq 2$. We then have the following result, which is used to derive the oracle inequality.
@@
Write $B_{T}=B\vee\left|\log(C_{T}/T)\right|$, it follows that $|g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})|\leq B_{T}$. The rest of the proof will be divided into six steps to enhance the readability.
@@
Let $\tilde{\mathcal{D}}_{T}=\{(\mathbf{\tilde{Z}}^{(t)},\mathbf{\tilde{y}}^{(t)})\}_{t=1}^{T}$ be i.i.d. random variables with same distribution as $(\mathbf{Z},\mathbf{y})$ and independent of $\mathcal{D}_{T}$. Consider the event
@@
where the second equation is due to the fact that $\mathbf{y}$ is one-hot encoded, and the second inequality follows from Lemma 25 in the following. Since for any $\mathbf{p}\in\mathcal{F}$, we have $p_{k}(\mathbf{Z})\geq C_{1}/T^{C_{2}}$ for some constants $C_{1}>0$ and $0<C_{2}<1$, and $C_{T}\leq C_{1}T^{1-C_{2}}$, it follows that
@@
Now we investigate the last probability in (21) condition on $\mathcal{D}_{T}$. Fixing $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}^{(t)}):t=1,\dots,T\}$, let $\delta>0$ and let $\mathcal{G}_{\delta}=\{g_{i}:i=1,\dots,M\}$ be a $\delta$-cover of $\mathcal{F}_{1}=\{g_{\mathbf{p}}:\mathbf{p}\in\mathcal{F}\}$ with respect to the norm $\|\cdot\|_{L_{1},\mathcal{D}_{T}}$ on the sample $\mathcal{D}_{T}$, where $M=N(\delta,\mathcal{F}_{1},\|\cdot\|_{L_{1},\mathcal{D}_{T}})$. Without loss of generalization, we assume that $\mathcal{G}_{\delta}\subset\mathcal{F}_{1}$. For a given $\mathbf{p}\in\mathcal{F}$, there exist $g_{i}\in\mathcal{F}_{1}$ such that
@@
This completes the proof of this Lemma.
The following result is the Lemma 3.7 in Bos and Schmidt-Hieber (2022), which is used in Step 1 of the proof of Lemma 24 with $m=2$. We call $(p_{1},\dots,p_{K})\in\mathbb{R}^{K}$ a probability vector if $\sum_{i=1}^{K}p_{i}=1$ and $p_{i}\geq 0$ for all $1\leq i\leq K$.
@@ Appendix B Proof of Theorem 9
Proof  By composing the neural networks established in Lemma 8 and Lemma 27, we construct a neural network $\mathbf{G}(\boldsymbol{\phi})=\big{(}G(\phi_{1}),\dots,G(\phi_{K})\big{)}$ such that, for any $\varepsilon\in(0,1)$,
@@
where $\boldsymbol{\phi}=(\phi_{1},\phi_{2},\dots,\phi_{K})^{\top}\in\mathcal{T}(M,D,H,W_{1},B_{1},S_{1},\gamma)$ and $\mathbf{G}\in\mathcal{F}(L,KW,KS)$. Consequently, for any $k=1,\dots,K$, we have
$\left\|e^{G(\phi_{k})}-p_{k}^{0}\right\|_{\infty}\leq\left\|e^{G(\phi_{k})}-\phi_{k}\right\|_{\infty}+\left\|\phi_{k}-p_{k}^{0}\right\|_{\infty}\leq\varepsilon+\varepsilon=2\varepsilon.$
@@
$\leq$ $2K\varepsilon+2\varepsilon=2\varepsilon(K+1).$
Next, we investigate the low bound of $\widetilde{q}_{k}(\mathbf{Z})$. According to Lemma  27, we have $e^{G(\phi_{k})}\geq\varepsilon$. Also note that
$\left\|\sum_{j=1}^{K}e^{G(\phi_{j}(\mathbf{Z}))}-1\right\|_{\infty}=\left\|\sum_{j=1}^{K}\{e^{G(\phi_{j}(\mathbf{Z}))}-p_{j}^{0}\}\right\|_{\infty}\leq 2K\varepsilon.$
Consequently,
$\widetilde{q}_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2K\varepsilon+1}\geq\frac{\varepsilon}{2K+1}\geq\frac{\varepsilon}{2(K+1)}.$
@@
The approximation capabilities of deep ReLU networks have been extensively studied, establishing sharp rates for various function classes (Yarotsky, 2017; DeVore et al., 2021; Elbrächter et al., 2021; Berner et al., 2021; Siegel, 2023; Schmidt-Hieber, 2020; Shen et al., 2019; Lu et al., 2021; Shen et al., 2022; Jiao et al., 2023). A key component in our subsequent analysis is the efficient approximation of the logarithm function by a scalar-valued neural network with controlled complexity. To this end, we present the following lemma, which directly adapts the construction from Theorem 4.1 of Bos and Schmidt-Hieber (2022).
Lemma 26 (Scalar approximation of $\log x$)
For any $\varepsilon\in(0,1)$ and $\beta>0$, there exists a scalar-valued ReLU neural network $G:[0,1]\to\mathbb{R}$ such that
$\left|e^{G(x)}-x\right|\leq\varepsilon,\quad\text{and}\quad G(x)\geq\log(\varepsilon),\quad\text{for all }x\in[0,1],$
with network depth $L$, width $W$, and size $S$ (total number of non-zero parameters) satisfying
$L\lesssim\log\left(\frac{1}{\varepsilon}\right),\quad W\lesssim\left(\frac{1}{\varepsilon}\right)^{\frac{1}{\beta}},\quad S\lesssim\left(\frac{1}{\varepsilon}\right)^{\frac{1}{\beta}}\log\left(\frac{1}{\varepsilon}\right).$
Lemma 27 (Vectorized approximation via coordinate-wise application)
Let $K\geq 1$, and let $G$ be a scalar-valued neural network with depth $L$, width $W$, and at most $S$ nonzero parameters. Define the vector-valued network
$\mathbf{G}(\boldsymbol{\phi}(\mathbf{Z})):=\bigl{(}G(\phi_{1}(\mathbf{Z})),\dots,G(\phi_{K}(\mathbf{Z}))\bigr{)}^{\top},$
where $\boldsymbol{\phi}(\mathbf{Z})=(\phi_{1}(\mathbf{Z}),\dots,\phi_{K}(\mathbf{Z}))^{\top}\in[0,1]^{K}$. Then $\mathbf{G}\in\mathcal{F}_{\mathrm{id}}(L,KW,KS)$, and for all $k\in[K]$ and $\mathbf{Z}\in[0,1]^{d}$,
$\left|e^{G(\phi_{k}(\mathbf{Z}))}-\phi_{k}(\mathbf{Z})\right|\leq\varepsilon,\quad\text{and}\quad G(\phi_{k}(\mathbf{Z}))\geq\log(\varepsilon).$
Proof  Although $\mathcal{F}_{\mathrm{id}}(L,W,S)$ as defined in (5) refers to vector-valued neural networks mapping $0,1]^{K}\to\mathbb{R}^{K}$, we temporarily allow the notation to refer to scalar-valued networks when the context is clear.
To prove the lemma, it suffices to construct a scalar-valued ReLU neural network $G:[0,1]\to\mathbb{R}$ such that
$\left|e^{G(x)}-x\right|\leq\varepsilon\quad\text{and}\quad G(x)\geq\log(\varepsilon)\quad\text{for all }x\in[0,1].$
Given such a function $G$, we define the vector-valued network $\mathbf{G}:[0,1]^{K}\to\mathbb{R}^{K}$ by
$\mathbf{G}(\mathbf{x}):=\bigl{(}G(x_{1}),G(x_{2}),\dots,G(x_{K})\bigr{)}^{\top}.$
By construction, for each $k\in[K]$, the coordinate function satisfies
$\left|e^{G(x_{k})}-x_{k}\right|\leq\varepsilon,\quad\text{and}\quad G(x_{k})\geq\log(\varepsilon),\quad\text{for all }x_{k}\in[0,1],$
so $\mathbf{G}$ satisfies the desired approximation and positivity conditions coordinate-wise. Since the same network $G$ is applied independently to each coordinate, the resulting network $\mathbf{G}$ lies in the class $\mathcal{F}_{\mathrm{id}}(L,KW,KS)$, completing the proof.
@@ Appendix C Proof of Theorem 11
To establish an upper bound for the risk $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$ in terms of the Kullback–Leibler (KL) divergence, we first provide two Lemmas that will be used in the proof of Theorem 11.
@@
Set
$g(\mathbf{z})=\frac{1}{p_{k}(\mathbf{z})}\mathbf{1}_{\{p_{k}(\mathbf{z})\geq h\}},$
thus
$\int_{\{p_{k}(\mathbf{z})\geq h\}}\frac{1}{p_{k}(\mathbf{z})}\,d\mathbb{P}_{\mathbf{Z}}(\mathbf{z})=\int_{0}^{\infty}\mathbb{P}_{\mathbf{Z}}\left(\frac{1}{p_{k}(\mathbf{Z})}\mathbb{I}_{\{p_{k}(\mathbf{Z})\geq h\}}\geq u\right)\,du.$
@@
$\frac{1}{p_{k}(\mathbf{Z})}\mathbf{1}_{\{p_{k}(\mathbf{Z})\geq h\}}\geq u>0\quad\Longrightarrow\quad h\leq p_{k}(\mathbf{Z})\leq\frac{1}{u},$
@@
Let $P$ and $Q$ be two probability measures satisfying $P\ll Q$ with the Radon-Nikodym derivative
$g(z)=\frac{dP}{dQ}(z).$
Then, the following inequality holds:
@@
Proof  Recall that $R(\mathbf{p}_{0},\mathbf{p})=\mathbb{E}_{\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\mathbf{p}(\mathbf{Z})}\right].$ Applying Lemma 29 to our setting yields
@@
$\|\mathbf{p}-\mathbf{p}_{0}\|_{\infty}\leq 2(K+1)\varepsilon,\quad\text{and}\quad\min_{1\leq k\leq K}\inf_{\mathbf{Z}\in[0,1]^{d}}p_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2(K+1)}.$
@@
$p_{k}(\mathbf{Z})\geq\max\left\{p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon,\;\frac{\varepsilon}{2(K+1)}\right\}.$
@@
$p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon\geq\frac{\varepsilon}{2(K+1)},$
@@
$p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon\geq p_{k}^{0}(\mathbf{Z})\left(1-\frac{2(K+1)}{2(K+1)+\frac{1}{2(K+1)}}\right)=\frac{p_{k}^{0}(\mathbf{Z})}{1+4(K+1)^{2}}.$
@@
$\begin{split}\frac{\left(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z})\right)^{2}}{p_{k}(\mathbf{Z})}&\leq 8(K+1)^{3}\varepsilon\cdot\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})\leq 2(K+1)\varepsilon+\frac{\varepsilon}{2(K+1)}\right\}\\ &+\frac{4(K+1)^{2}\left[1+4(K+1)^{2}\right]\varepsilon^{2}}{p_{k}^{0}(\mathbf{Z})}\cdot\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})>2(K+1)\varepsilon+\frac{\varepsilon}{2(K+1)}\right\}.\end{split}$
@@
$\mathbb{E}_{\mathbf{Z}}\left[\frac{\left(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z})\right)^{2}}{p_{k}(\mathbf{Z})}\right]\leq 8(K+1)^{3}\varepsilon\,\mathbb{P}\left\{p_{k}^{0}(\mathbf{Z})\leq 2(K+1)\varepsilon+\frac{\varepsilon}{2(K+1)}\right\}$
$+4(K+1)^{2}\left[1+4(K+1)^{2}\right]\varepsilon^{2}\,\mathbb{E}_{\mathbf{Z}}\left[\frac{\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})>2(K+1)\varepsilon+\frac{\varepsilon}{2(K+1)}\right\}}{p_{k}^{0}(\mathbf{Z})}\right].$
@@
$\leq$ $C\,\varepsilon^{1+(\alpha\wedge 1)}\,8(K+1)^{3}\left(\frac{4(K+1)^{2}+1}{2(K+1)}\right)^{\alpha\wedge 1}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right),$
where $\alpha\wedge\beta=\min\{\alpha,\beta\}$. Also note that $0\leq\alpha\leq 1$ and $\frac{4(K+1)^{2}+1}{2(K+1)}\leq 3(K+1)$. Consequently,
@@
$\leq C\varepsilon^{1+(\alpha\wedge 1)}K(K+1)^{3}\left(\frac{4(K+1)^{2}+1}{2(K+1)}\right)^{\alpha\wedge 1}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\!\left(\frac{1}{\varepsilon}\right)\right)$
$\leq C\varepsilon^{1+(\alpha\wedge 1)}K(K+1)^{3+(\alpha\wedge 1)}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right)$
$=C\varepsilon^{1+\alpha}K(K+1)^{3+\alpha}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right)$
@@
This concludes the proof of Theorem  11.

### v2

@@ Appendix A Proof of Theorem 7
  • The expected excess risk defined as:
@@
  • The expected empirical risk gap is defined as:
@@ A.1 Supporting Lemmas
$\bigg{|}\mathbb{E}_{\mathcal{D}_{T}}\left[y_{k}^{(t)}\left(B\wedge\log\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{\widehat{p}_{k}(\mathbf{Z}^{(t)})}\right)\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})<\frac{C_{T}}{T}\right)\right]\bigg{|}\leq\frac{C_{T}\{\log(T/C_{T})+B\}}{T},$
@@
$\bigg{|}\mathbb{E}_{\mathcal{D}_{T}}\left[y_{k}^{(t)}\left(B\wedge\log\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{\widehat{p}_{k}(\mathbf{Z}^{(t)})}\right)\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})<\frac{C_{T}}{T}\right)\right]\bigg{|}$
@@
This completes the proof of Lemma 24.
@@
Recall that $\mathbf{p}(\mathbf{Z})=(p_{1}(\mathbf{Z}),\dots,p_{K}(\mathbf{Z}))^{\top}$and $g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})=\mathbf{y}^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\mathbf{p}(\mathbf{Z})}\right)_{\geq C_{T}/T}$, where $B\geq 2$. We then have the following result, which is used to derive the oracle inequality.
@@
Write $B_{T}=B\vee\left|\log(C_{T}/T)\right|$, and it follows that $|g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})|\leq B_{T}$. The rest of the proof will be divided into six steps to enhance the readability.
@@
Let $\tilde{\mathcal{D}}_{T}=\{(\mathbf{\tilde{Z}}^{(t)},\mathbf{\tilde{y}}^{(t)})\}_{t=1}^{T}$ be i.i.d. random variables with the same distribution as $(\mathbf{Z},\mathbf{y})$ and independent of $\mathcal{D}_{T}$. Consider the event
@@
where the second equation is due to the fact that $\mathbf{y}$ is one-hot label vector, and the second inequality follows from Lemma 27 in the following. Since for any $\mathbf{p}\in\mathcal{F}$, we have $p_{k}(\mathbf{Z})\geq C_{1}/T^{C_{2}}$ for some constants $C_{1}>0$ and $0<C_{2}<1$, and $C_{T}\leq C_{1}T^{1-C_{2}}$, it follows that
@@
Now we study the last probability in (21), conditioned on $\mathcal{D}_{T}$. Fix $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}^{(t)}):t=1,\dots,T\}$, let $\delta>0$, and define $\mathcal{G}_{\delta}=\{g_{i}:i=1,\dots,M\}$ as a $\delta$-cover of $\mathcal{F}_{1}=\{g_{\mathbf{p}}:\mathbf{p}\in\mathcal{F}\}$ with respect to $\|\cdot\|_{L_{1},\mathcal{D}_{T}}$, where $M=N(\delta,\mathcal{F}_{1},\|\cdot\|_{L_{1},\mathcal{D}_{T}})$.
Without loss of generalization, we assume that $\mathcal{G}_{\delta}\subset\mathcal{F}_{1}$. For a given $\mathbf{p}\in\mathcal{F}$, there exist $g_{i}\in\mathcal{F}_{1}$ such that
@@
This completes the proof of Lemma 26.
The following lemma, adapted from Lemma 3.7 in Bos and Schmidt-Hieber (2022), is used in Step 1 of the proof of Lemma 26 with $m=2$. We call $(p_{1},\dots,p_{K})\in\mathbb{R}^{K}$ a probability vector if $\sum_{i=1}^{K}p_{i}=1$ and $p_{i}\geq 0$ for all $1\leq i\leq K$.
@@ Appendix B Proof of Theorem 9
The following lemma is a key component in the proof of Theorem 9. It provides an efficient approximation of the logarithm function by a scalar-valued neural network with controlled complexity. This result is adapted from Theorem 4.1 of Bos and Schmidt-Hieber (2022), with some modifications to the notations.
Lemma 28 (Theorem 4.1 of Bos and Schmidt-Hieber (2022))
For any $\varepsilon\in(0,1)$ and $\beta>0$, there exists a scalar-valued ReLU neural network $G:[0,1]\to\mathbb{R}$ with the depth $L$, width $W$, and size $S$ (the total number of nonzero parameters) satisfying
$L\lesssim\log\left(\frac{1}{\varepsilon}\right),\quad W\lesssim\left(\frac{1}{\varepsilon}\right)^{\frac{1}{\beta}},\quad S\lesssim\left(\frac{1}{\varepsilon}\right)^{\frac{1}{\beta}}\log\left(\frac{1}{\varepsilon}\right),$
such that
$\left|e^{G(x)}-x\right|\leq\varepsilon,\quad\text{and}\quad G(x)\geq\log(\varepsilon),\quad\text{for all }x\in[0,1].$
Remark 29
Note that the neural network $G$ constructed in Lemma 28 also satisfies $G(x)\leq\log(x+\varepsilon)\leq\log(2)$. This upper bound of $G(x)$ will be used in the proof of Theorem 9.
Proof [Proof of Theorem 9] By composing the neural networks established in Lemma 8 and Lemma 28, we construct a neural network $\mathbf{G}(\boldsymbol{\phi})=\big{(}G(\phi_{1}),\dots,G(\phi_{K})\big{)}$ such that, for any $\varepsilon\in(0,1)$,
@@
where $\boldsymbol{\phi}=(\phi_{1},\phi_{2},\dots,\phi_{K})^{\top}\in\mathcal{T}(M,D,H,W_{1},B_{1},S_{1},\gamma)$ and $G\in\mathcal{F}_{\mathrm{id}}(L,W,S)$ are provided in Lemma 8 and Lemma 28, respectively. Here we temporarily allow the notation $\mathcal{F}_{\mathrm{id}}(L,W,S)$ to refer to scalar-valued networks with depth $L$, width $W$, and size $S$. It is readily seen that $\mathbf{G}\in\mathcal{F}_{\mathrm{id}}(L,KW,KS)$. Consequently, for any $k=1,\dots,K$, we have
$\left\|e^{G(\phi_{k})}-p_{k}^{0}\right\|_{\infty}\leq\left\|e^{G(\phi_{k})}-\phi_{k}\right\|_{\infty}+\left\|\phi_{k}-p_{k}^{0}\right\|_{\infty}\leq 2\varepsilon.$
@@
$\leq$ $2K\varepsilon+2\varepsilon$
$=$ $2\varepsilon(K+1).$
Next, we investigate the low bound of $\widetilde{q}_{k}(\mathbf{Z})$. According to Lemma  28 and Remark 29, we have $\log(\varepsilon)\leq G(\phi_{k})\leq\log(2)$. Consequently,
$\widetilde{q}_{k}(\mathbf{Z})=\frac{e^{G(\phi_{k}(\mathbf{Z}))}}{\sum_{j=1}^{K}e^{G(\phi_{j}(\mathbf{Z}))}}\geq\frac{\varepsilon}{2K}.$
@@
Remark 30
According to the proof of Theorem 4.1 of Bos and Schmidt-Hieber (2022), for any ReLU neural network $G_{0}$ and $\varepsilon\in(0,1)$, we can always construct a network $G(x)$ by adding additional ReLU layers to $G_{0}$ such that $G(x)=\{G_{0}(x)\vee\log(\varepsilon)\}\wedge\log(2)$, where $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$. If $|e^{G_{0}(x)}-x|\leq\varepsilon$ for all $x\in[0,1]$, then $|e^{G(x)}-x|\leq\varepsilon$ for all $x\in[0,1]$. This construction also ensures $\log(\varepsilon)\leq G(x)\leq\log(2)$. Therefore, by applying this construction to the MLP architecture in the network of $\mathcal{F}$ and adapting the arguments in proof of Theorem 9, we can ensure that for any network $\mathbf{p}=(p_{1},\dots,p_{K})\in\mathcal{F}$,
$p_{k}\geq\frac{\varepsilon}{2K},\quad k=1,\dots,K,$
where $\mathcal{F}=\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,W,S)$ is defined in (3.4).
@@ Appendix C Proof of Theorem 10
To establish an upper bound for the approximation error $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$ in terms of the Kullback-Leibler (KL) divergence, we first provide two Lemmas that will be used in the proof of Theorem 10.
@@
Write $g(\mathbf{z})=\frac{1}{p_{k}(\mathbf{z})}\mathbb{I}(p_{k}(\mathbf{z})\geq h)$, it follows that
$\int_{\{p_{k}(\mathbf{z})\geq h\}}\frac{1}{p_{k}(\mathbf{z})}\,d\mathbb{P}_{\mathbf{Z}}(\mathbf{z})=\int_{0}^{\infty}\mathbb{P}_{\mathbf{Z}}\left(\frac{1}{p_{k}(\mathbf{Z})}\mathbb{I}(p_{k}(\mathbf{z})\geq h)\geq u\right)\,du.$
@@
$\frac{1}{p_{k}(\mathbf{Z})}\mathbb{I}(p_{k}(\mathbf{z})\geq h)\geq u>0\quad\Longrightarrow\quad h\leq p_{k}(\mathbf{Z})\leq\frac{1}{u},$
@@
Let $P$ and $Q$ be two probability measures satisfying $P\ll Q$ with the Radon-Nikodym derivative $g(z)=\frac{dP}{dQ}(z)$. Then, the following inequality holds:
@@
Proof [Proof of Theorem 10] Recall that $R(\mathbf{p}_{0},\mathbf{p})=\mathbb{E}_{\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\mathbf{p}(\mathbf{Z})}\right].$ Applying Lemma 32 to our setting yields
@@
$\|\mathbf{p}-\mathbf{p}_{0}\|_{\infty}\leq 2(K+1)\varepsilon,\quad\text{and}\quad\min_{1\leq k\leq K}\inf_{\mathbf{Z}\in[0,1]^{d}}p_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2K}.$
@@
$p_{k}(\mathbf{Z})\geq\max\left\{p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon,\;\frac{\varepsilon}{2K}\right\}.$
@@
$p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon\geq\frac{\varepsilon}{2K},$
@@
$p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon\geq p_{k}^{0}(\mathbf{Z})\left(1-\frac{2(K+1)}{2(K+1)+\frac{1}{2K}}\right)=\frac{p_{k}^{0}(\mathbf{Z})}{1+4K(K+1)}.$
@@
$\begin{split}\frac{\left(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z})\right)^{2}}{p_{k}(\mathbf{Z})}&\leq 8K(K+1)^{2}\varepsilon\cdot\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})\leq 2(K+1)\varepsilon+\frac{\varepsilon}{2K}\right\}\\ &\quad+\frac{4(K+1)^{2}\left[1+4K(K+1)\right]\varepsilon^{2}}{p_{k}^{0}(\mathbf{Z})}\cdot\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})>2(K+1)\varepsilon+\frac{\varepsilon}{2K}\right\}.\end{split}$
@@
$\mathbb{E}_{\mathbf{Z}}\left[\frac{\left(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z})\right)^{2}}{p_{k}(\mathbf{Z})}\right]\leq 8K(K+1)^{2}\varepsilon\,\mathbb{P}\left\{p_{k}^{0}(\mathbf{Z})\leq 2(K+1)\varepsilon+\frac{\varepsilon}{2K}\right\}$
$\quad+4(K+1)^{2}\left[1+4K(K+1)\right]\varepsilon^{2}\,\mathbb{E}_{\mathbf{Z}}\left[\frac{\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})>2(K+1)\varepsilon+\frac{\varepsilon}{2K}\right\}}{p_{k}^{0}(\mathbf{Z})}\right].$
@@
$\leq$ $C\,\varepsilon^{1+(\alpha\wedge 1)}\,8K(K+1)^{2}\left(\frac{4K(K+1)+1}{2K}\right)^{\alpha\wedge 1}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right),$
where $\alpha\wedge\beta=\min\{\alpha,\beta\}$. Also note that $0\leq\alpha\leq 1$ and $\frac{4K(K+1)+1}{2K}\leq 3(K+1)$. Consequently,
@@
$\leq C\varepsilon^{1+(\alpha\wedge 1)}K^{2}(K+1)^{2}\left(\frac{4K(K+1)+1}{2K}\right)^{\alpha\wedge 1}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\!\left(\frac{1}{\varepsilon}\right)\right)$
$\leq C\varepsilon^{1+(\alpha\wedge 1)}K^{2}(K+1)^{2+(\alpha\wedge 1)}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right)$
$=C\varepsilon^{1+\alpha}K^{2}(K+1)^{2+\alpha}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right)$
@@
This concludes the proof of Theorem 10.